An overnight rainstorm has caused a major roadblock. Three massive rocks of mass m1=598kg, m2=796kg and m3=311kg have blocked a busy road. The city calls a local contractor to use a bulldozer to clear the road. The bulldozer applies a constant force to m1 to slide the rocks off the road. Assuming the road is a flat frictionless surface and the rocks are all in contact, what force, FA must be applied to m1 to slowly accelerate the group of rocks from the road at 0.100m/s2. Use the value found above fro FA to find the force f12 exerted by the first rock of mass 598kg on the middle rock of 769kg
Answer #1

total mass = 598+769+311 kg = 1678 kg

acceleration = 0.1 m/s^2

so Fa = ma = 1678*0.1 = 167.8 N


m2 and m3 moves with force f12

so f12 = 0.1*(769+311) = 108 N

so F12 = 108 N
Answer #2

Fa = 167.8 N

F12 = 108 N